Time Frequency Brown Bag Seminar

Thursday May 21, 1998

12:00 noon

EQuad E415

Speaker: Yao Wang Associate Professor, Polytechnic University, Brooklyn, NY Visiting Research Staff, Princeton University

Title: Multiple Description Coding Using Pairwise Correlating Transforms

Abstract:

This talk considers the problem of signal transmission over unreliable networks in which more than one paths exist between the source and destination and that any path could experience failures. One way to enhance the reliability of such a communication system is by using multiple description coding (MDC) at the source coder. With MDC, two (or more) coded bit-streams (referred to as descriptions) of the same signal are generated. The coder is designed so that a high quality signal reproduction can be achieved with both descriptions, but the quality of the decoded signal is acceptable even with only one description. The MDC objective is realized by forcing dependencies between the coded descriptions, allowing either descriptio to be estimated from the other when one is lost. This dependency however will reduce the source coding efficiency. The challenge in designing a MDC coder is how to achieve a good trade-off between the coding efficiency and the reconstruction quality from a single description.

In this talk, we describe a MDC coder that is built on the standard transform coding framework. Unlike standard transform coding by which a linear transform is used to decorrelate coefficients to be coded and transmitted, we use the transform to force some structured correlation among coefficients. We will present analysis results obtained for the Gaussian source model, using a new concept of redundancy rate-distortion function. We will also discuss the applications of the MDC coder for image and video transmission over wireless and packet networks and show some preliminary simulation results with images.

Joint work with Michael Orchard, Amy Reibman and Vinay Vaishampayan